# TritonRL: Training LLMs to Think and Code Triton Without Cheating

### v1

@@ Abstract
With the rapid evolution of large language models (LLMs), the demand for automated, high-performance system kernels has emerged as a key enabler for accelerating development and deployment. We introduce TritonRL, a domain-specialized LLM for Triton kernel generation, trained with a novel training framework that enables robust and automated kernel synthesis. Unlike general-purpose programming languages, Triton kernel generation faces unique challenges due to data scarcity and incomplete evaluation criteria, vulnerable to reward hacking. Our approach addresses these challenges end-to-end by distilling Triton-specific knowledge through supervised fine-tuning on curated datasets, and further improving code quality via reinforcement learning (RL) with robust, verifiable rewards and hierarchical reward assignment. Our RL framework robustly detects reward hacking and guides both reasoning traces and code tokens through fine-grained verification and hierarchical reward decomposition, enabling the model to generate high-quality Triton kernels that can truly replace existing modules. With robust and fine-grained evaluation, our experiments on KernelBench demonstrate that TritonRL achieves state-of-the-art correctness and speedup, surpassing all other Triton-specific models and underscoring the effectiveness of our RL-based training paradigm.
@@ 1 Introduction
The exponential growth in demand for GPU computing resources has driven the need for highly optimized GPU kernels that improve computational efficiency, yet with the emergence of numerous GPU variants featuring diverse hardware specifications and the corresponding variety of optimization kernels required for each, developing optimized kernels has become an extremely time-consuming and challenging task. In response to this need, there is growing interest in leveraging large language models (LLMs) for automated kernel generation. While there have been attempts introducing inference frameworks that utilize general-purpose models, such as OpenAI models and DeepSeek, for generating kernels (Ouyang et al., 2025; Lange et al., 2025; Li et al., 2025a; NVIDIA Developer Blog, 2025; Muhamed et al., 2023), they often struggle with even basic kernel implementations, thereby highlighting the critical need for domain-specific models specifically tailored for kernel synthesis.
As the need for specialized models for kernel generation has emerged, several works have focused on fine-tuning LLMs for CUDA or Triton. In the CUDA domain, recent RL-based approaches include Kevin-32B Baronio et al. (2025), which progressively improves kernels using execution feedback as reward signals, and CUDA-L1, which applies contrastive RL to DeepSeek-V3 Li et al. (2025c). While these large models (32B-671B parameters) achieve strong CUDA performance, their training costs remain prohibitively expensive. To address these limitations, researchers have proposed specialized 8B Triton models, including KernelLLM Fisches et al. (2025) (supervised training on torch compiler-generated code) and AutoTriton Li et al. (2025b), fine-tuned via LLM distillation and RL using execution feedback. Though these smaller models outperform their base models, significant room remains for improving efficiency and accuracy compared to larger counterparts.
@@
In this paper, we present TritonRL, an 8B-scale LLM specialized for Triton programming that achieves state-of-the-art performance in both correctness and runtime speedup, while effectively mitigating reward hacking. To enable high-quality Triton kernel generation with small models (up to 8B), we design a training pipeline with the following key contributions:
  • Simplified dataset curation with distillation: Instead of large-scale web crawling, we build on the curated KernelBook problems. Their solutions are refined and augmented through DeepSeek-R1 distillation (Guo et al., 2025), providing high-quality supervision for SFT of our base model Qwen3-8B (Team, 2025).
  • Fine-grained and robust verification: We incorporate enhanced rule-based checks (e.g., nn.Module) together with LLM-based judges (e.g., Qwen3-235B-Instruct) to construct verifiable rewards. This enables reliable diagnosis across commercial and open-source models, while preventing reward hacking that arises from naive syntax-only verification.
  • Hierarchical reward decomposition with data mixing: Our RL stage decomposes rewards into multiple dimensions (e.g., correctness, efficiency, style) and applies token-level credit assignment. Combined with strategic data mixing across SFT and RL, this yields better kernel quality, generalization, and robustness.
  • Comprehensive evaluation and open-sourcing: Through rigorous validity analysis that filters out syntactically or functionally invalid code, we reveal true performance differences among models. Ablation studies further confirm the effectiveness of our hierarchical reward design and data mixing. At the 8B scale, TritonRL surpasses existing Triton-specific LLMs, including KernelLLM (Fisches et al., 2025) and AutoTriton (Li et al., 2025b). We release our datasets and codes, including training and evaluation frameworks, to ensure reproducibility and foster future research. [^1] Codes and datasets will be available soon.
@@ 2 TritonRL
Figure 1: TritonRL components and workflow.
In this section, we present TritonRL, a specialized model designed for Triton programming. Our objective is to develop a model that can generate Triton code that is both correct and highly optimized for speed, outperforming the reference implementation. To achieve this, we adopt a two-stage training strategy: we first apply supervised fine-tuning (SFT) to instill fundamental Triton syntax and kernel optimization skills, followed by reinforcement learning (RL) with fine-grained verifiable rewards to further refine the model for correctness and efficiency. We will first detail the SFT procedure, and subsequently present the design of the reinforcement learning framework.
2.1 Triton Knowledge Distillation via Supervised Fine-Tuning
Recent work (Fisches et al., 2025; Li et al., 2025b) has demonstrated that large language models (LLMs) exhibit weak Triton programming ability at the 8B scale, struggling both with syntax and with performance-oriented design patterns. Effective dataset curation is therefore essential, not only to expose Triton-specific primitives and coding structures but also to preserve the reasoning traces that guide kernel optimization. To address this, our pipeline follows three key steps: (i) data augmentation, (ii) synthesis of reasoning traces paired with corresponding code, and (iii) construction of high-quality training pairs for supervised fine-tuning.
  1. (i)
    Data augmentation: We start from the problem sets in KernelBook (Paliskara & Saroufim, 2025), which provides curated pairs of PyTorch programs and equivalent Triton kernels. To enrich this dataset, we augment the tasks with additional variations (e.g., diverse input shapes), thereby exposing the model to broader performance scenarios.
  2. (ii)
    Data synthesis: To obtain diverse reasoning traces that guide correct and efficient Triton generation, we employ DeepSeek-R1 (Guo et al., 2025) to jointly synthesize reasoning steps and Triton implementations. For each task wrapped with instruction and Pytorch reference, multiple candidate kernels are collected, each paired with an explicit reasoning trace. This yields a dataset of $\mathcal{D}=\big\{(q,o_{i})\big\}=\{(\texttt{task query, Triton code with CoT})\}$. See concrete template in Appendix F.5). Rather than directly using the Triton code from KernelBook as the target, we instead utilize Triton implementations generated by DeepSeek-R1. This choice ensures higher quality supervision, as KernelBook’s original Triton code frequently relies on high-level PyTorch modules (see Appendix F.2 for examples).
  3. (iii)
    Supervised fine-tuning: In the SFT stage, the model is trained to produce valid Triton code as well as reproduce the associated reasoning traces from the instruction. This distillation process transfers essential Triton programming patterns while reinforcing reasoning ability.
2.2 Reinforcement Learning with Hierarchical Reward Decomposition
While supervised fine-tuning (SFT) equips the base model with basic Triton syntax and kernel optimization abilities, the resulting code may still contain errors or lack efficiency. To further improve the quality of Triton code generation of TritonRL, we train the model via reinforcement learning (RL) with verifiable rewards, which incentivize model to generate more correct and efficient Triton code yielding higher rewards. In RL, designing effective reward feedback is essential since crude reward designs not perfectly aligned with original objectives of tasks often lead to reward hacking, guiding models to exploit loopholes. To address this, we first introduce robust and fine-grained verifiers that rigorously assess the quality of Triton code in diverse aspects, forming the basis for constructing reward functions. Building on these verifiers, we present a GRPO-based RL framework with hierarchical reward decomposition that provides targeted feedback for reasoning traces and Triton code, thereby improving the correctness and efficiency of generated Triton kernels.
@@ 2.2 Reinforcement Learning with Hierarchical Reward Decomposition
Figure 2: Illustration of the flow of our robust verifier incorporating syntax and functionality checkers and the examples of invalid Triton codes. (a) invalid syntax: the code lacks any Triton blocks and consists solely of PyTorch code. (b) invalid functionality: the code include dummy Triton code that just copies data without meaningful operation delegating core operation (matrix multiplication) to PyTorch modules (torch.matmul).
Fine-Grained Verification for High-Quality Triton Code.
We denote $i$-th generated output sample $o_{i}$ for a given prompt or task $q$. Recall that $q$ and $o_{i}$ include Pytorch reference $q^{\text{ref}}$ and Triton code $o_{i}^{\text{code}}$ that is executable on some proper input $x$, i.e., $q^{\text{ref}}(x),o^{\text{code}}(x)$. We introduce fine-grained verifiers $v$ that comprehensively evaluate different aspects of code quality as follows.
  • Robust verifier: Sequential verifier to check output is a valid Triton code without non-nonsensical hacking. See Figure 2 for an illustration
@@
    • func: A binary functionality verifier to detect whether $o^{\text{code}}$ constitutes a valid Triton kernel. Syntax checks alone are insufficient, since models may output code that superficially passes verification but defers computations to high-level PyTorch modules (e.g., torch.nn, @) or hardcodes constants, as in Figure 2 (b). To address this, we combine a rule-based linter—which ensures Triton kernels are invoked and flags reliance on PyTorch modules—with an LLM-based judge that evaluates semantic correctness against task specifications.
@@
  • compiled: A binary verifier that checks whether the generated Triton code can be successfully compiled without errors.
@@
  • correct: A binary verifier that evaluates whether the generated Triton code produces correct outputs by compiling and comparing its results against those of the reference PyTorch code using provided test input $x$.
@@
    $\texttt{correct}(q,o)\;=\;\texttt{compile}\!\left(o^{\text{code}}\right)\cdot\mathbbm{1}\!\left[\,o^{\text{code}}(x)==q^{\text{ref}}(x)\,\right]$
@@
  • speedup: A scalar score that quantifies the execution time improvement of the generated Triton code relative to the reference PyTorch implementation. For a prompt $q$ with reference Pytorch code $q^{\text{ref}}$ and corresponding triton code $o^{\text{code}}$ generated, speed-up is defined as, given test input $x$,
@@
    $\texttt{speedup}(q,o)=\frac{\tau(o^{\text{code}},x)}{\tau(q^{\text{ref}},x)}\cdot\texttt{correct}(q,o),$
@@
    where $\tau(\cdot,x)$ measure the runtimes of given code and input $x$.
For notational simplicity, for a given prompt $q$ and output sample $o_{i}$, we define $v(q,o_{i})$ as $v_{i}$ any verification function $v$ throughout the paper.
While prior works (Li et al., 2025b; Baronio et al., 2025) addressed reward hacking with rule-based linters, such methods remain vulnerable to loopholes. Our verifier combines rule-based and LLM-based checks to capture both syntactic and semantic errors, offering stronger guidance during training (see Appendix F.4 for examples and Section 3 for evaluation). Building on these fine-grained verifiers, we develop an RL framework that delivers targeted feedback to both reasoning traces and Triton code, improving kernel correctness and efficiency.
Hierarchical Reward Decomposition.
Training LLMs with long reasoning traces remains challenging because providing appropriate feedback across lengthy responses is difficult. When a single final reward is uniformly applied to all tokens, it fails to distinguish between those that meaningfully contribute to correctness or efficiency and those that do not. This issue is particularly pronounced in kernel code generation, where reasoning traces often outline complex optimization strategies for GPU operations while the subsequent code implements these plans. Even if the reasoning trace proposes a promising optimization, errors in the Triton implementation may result in the entire response being penalized. Thus, a single reward signal fails to provide appropriate feedback to both reasoning and solution (Qu et al., 2025), conflating high-quality reasoning with poor execution and preventing effective learning of good optimization strategies.
To address this, we propose a GRPO with hierarchical reward decomposition for Triton code generation. Specifically, Triton code generation $o_{i}$ can be viewed as two-level hierarchy action pairs,
  • $o_{i}^{\text{plan}}$: CoT reasoning traces correspond to high-level planning actions, providing abstract kernel optimization strategies, such as tiling or shared memory.
  • $o_{i}^{\text{code}}$: final Triton code answers correspond to low-level coding actions that execute the plan given by the previous reasoning traces.
The key idea is to assign different reward credit for different class of output tokens between plan and code as illustrated in Figure 3. By jointly optimizing rewards for both levels, we can train the model to better align its reasoning with the desired code output as follows:

### v2

@@ Abstract
The rapid evolution of Large Language Models (LLMs) has driven a growing demand for automated, high-performance system kernels to accelerate machine learning workloads. We introduce TritonRL, a domain-specialized 8B-scale LLM for Triton programming, trained via a novel reinforcement learning (RL) framework. While Triton synthesis faces unique challenges, including data scarcity and a high susceptibility to reward hacking, our approach enables robust kernel generation through two primary innovations. First, we implement a multi-layered verification system that provides high-fidelity reward signals, ensuring that generated kernels are both syntactically and functionally valid. Second, we propose Hierarchical Reward Decomposition (HRD), which decouples reinforcement for high-level reasoning and low-level implementation to resolve the credit assignment problem in long-sequence generation. Comprehensive evaluations on KernelBench demonstrate that TritonRL achieves state-of-the-art correctness and runtime speedup, outperforming concurrent Triton-specific models and matching the performance of frontier models with over 100B parameters. Our results highlight the effectiveness of hardware-aware RL paradigms in specialized domain adaptation.
Machine Learning, ICML
@@ 1 Introduction
The exponential growth in demand for GPU computing resources has driven the need for highly optimized GPU kernels that improve computational efficiency, yet with the emergence of numerous GPU variants featuring diverse hardware specifications and the corresponding variety of optimization kernels required for each, developing optimized kernels has become an extremely time-consuming and challenging task. In response to this need, there is growing interest in leveraging large language models (LLMs) for automated kernel generation. While there have been attempts introducing inference frameworks that utilize general-purpose models, such as OpenAI models and DeepSeek, for generating kernels (Ouyang et al., 2025; Lange et al., 2025; NVIDIA Developer Blog, 2025; Muhamed et al., 2023), they often struggle with even basic kernel implementations, thereby highlighting the critical need for domain-specific models specifically tailored for kernel synthesis.
As the need for specialized models for kernel generation has emerged, several works have focused on fine-tuning LLMs for CUDA or Triton. In the CUDA domain, recent RL-based approaches include Kevin-32B Baronio et al. (2025), which progressively improves kernels using execution feedback as reward signals, and CUDA-L1, which applies contrastive RL to DeepSeek-V3 Li et al. (2025c). While these large models (32B-671B parameters) achieve strong CUDA performance, their training costs remain prohibitively expensive. Other research has focused on Triton, a domain-specific language that provides a higher-level abstraction than CUDA for writing efficient GPU kernels. Unlike CUDA, which benefits from decades of development and abundant training data, Triton is newer and has fewer high-quality kernel examples, making it more difficult to train specialized models. Recent works have trained LLMs for Triton kernel generation via supervised fine-tuning (SFT) on torch compiler–generated code (Fisches et al., 2025) or via LLM distillation followed by RL with execution feedback (Li et al., 2025b), typically at the 8B parameter scale. Although these smaller models improve over their base models, there remains substantial room to improve efficiency and correctness.
@@
In this paper, we present TritonRL, an 8B-scale LLM specialized for Triton programming that achieves state-of-the-art performance in both functional correctness and execution speed. Our training pipeline enables 8B-scale models to robustly surpass much larger frontiers through the following contributions:
  • Robust Verification Framework: We introduce a multi-layered verification system designed specifically to mitigate Triton-specific reward hacking. By identifying failure modes where models bypass core custom kernel logic, we implement fine-grained verifiers, utilizing both rule-based heuristics and LLM-based judges, that ensure reward signals are grounded in genuine functional and syntactic integrity.
  • Hierarchical Reward Decomposition (HRD): We propose a novel RL optimization objective that addresses the credit assignment problem in long-sequence generation. By decoupling rewards for high-level reasoning (planning) and low-level implementation (coding), we provide targeted reinforcement for strategy and execution independently. This approach outperforms uniform sequence-level rewards in both accuracy and speedup.
  • Comprehensive Evaluation and State-of-the-Art Performance: We provide a rigorous evaluation of open-source and proprietary models using our verification framework, revealing critical performance gaps in functional validity. Our results demonstrate that TritonRL sets a new state-of-the-art for 8B-scale models and performs comparably to or better than frontier models with over 100100B parameters. [^1] Codes and datasets will be available soon.
1.1 Related Work
Large Language Models for Kernel Generation
The demand for GPU efficiency has spurred interest in using large language models (LLMs) for automated kernel generation in CUDA and Triton (Ouyang et al., 2025; Li et al., 2025a; NVIDIA Developer Blog, 2025). Building on this, benchmarks like KernelBench (Ouyang et al., 2025) and TritonBench (Li et al., 2025a) have been introduced to evaluate LLMs on kernel optimization. Although LLMs show promise in general-purpose programming, they often achieve low success rates on specialized GPU programming tasks (Ouyang et al., 2025), highlighting the need for domain-specific models tailored to kernel synthesis.
Recent research focuses on domain-specific fine-tuning. For CUDA, models like Kevin-32B (Baronio et al., 2025) and CUDA-L1 (Li et al., 2025c) utilize multi-turn or contrastive RL to optimize performance. In the Triton domain, KernelLLM (Fisches et al., 2025) performs SFT on the KernelBook dataset (Paliskara and Saroufim, 2025), while AutoTriton (Li et al., 2025b) applies SFT and RL with verifiable rewards based on correctness. Both KernelLLM and AutoTriton are concurrent works developed alongside our work, and we provide a detailed performance comparison in Section 3.2.
Reinforcement Learning with Verifiable Rewards
Reinforcement Learning with Verifiable Rewards (RLVR) has become a standard for training LLMs in objective domains like mathematics and coding (Shao et al., 2024; Guo et al., 2025; Li et al., 2022). In these frameworks, the accuracy of the reward signal is critical to prevent reward hacking (Skalse et al., 2022), where models exploit test suites without genuinely solving the problem (Sharma et al., 2024; Gao et al., 2024). In kernel generation, reward hacking typically involves models using high-level PyTorch modules to bypass custom kernel implementation while still passing unit tests. While some recent efforts (Li et al., 2025b; Baronio et al., 2025) have introduced syntax checks, these checks remain insufficient to detect subtler forms of reward hacking such as computation delegation, leaving a critical gap in reward fidelity.
Figure 1: RL training workflow of TritonRL. The framework illustrates how kernel generation flows through fine-grained verification layers and hierarchical reward decomposition, providing robust and targeted feedback separately to reasoning traces (plan tokens) and to Triton code (code tokens) conditioned on those plans.
@@ 2 TritonRL
In this section, we introduce TritonRL, a comprehensive reinforcement learning (RL) framework designed to optimize language models for high performance Triton kernel generation. Our approach addresses the dual challenge of achieving functional correctness and hardware specific efficiency, optimizing for execution latency on target accelerators. To this end, our methodology is built upon three core components: (1) base model priming and task curation, (2) reward formulation with robust verification, and (3) policy optimization with hierarchical reward decomposition.
2.1 Base Model Priming and Task Curation
For a model to effectively optimize Triton kernels via RL, it must first possess a foundational understanding of Triton syntax and be exposed to a high quality task distribution during RL. We address these requirements through two efforts: (1) base model priming via knowledge distillation to instill fundamental programming skills and (2) task curation via data augmentation and difficulty aware data mixing to construct an optimal training distribution for RL.
2.1.1 Triton Knowledge Distillation via SFT
Recent studies indicate that pretrained language models, particularly at the 8B scale, exhibit limited proficiency in Triton programming, often struggling with low level syntax and hardware specific optimization (Fisches et al., 2025; Li et al., 2025b). To bridge the gap, we perform supervised fine-tuning (SFT) on a base model to distill expertise from frontier models into a more efficient base. In this paper, we utilize Qwen3-8B (Team, 2025) as our base model.
Our primary data source is KernelBook (Paliskara and Saroufim, 2025), which contains a rich collection of Triton kernel generation tasks paired with reference PyTorch implementations. We define $\mathcal{G}_{KB}=\{g_{i}\}_{i\in\{1,\dots,11K\}}$ as a set of 11K tasks sampled from KernelBook. For each task query $q_{i}=q(g_{i})$, wrapped with an instruction to generate a Triton kernel for task $g_{i}$, we prompt a teacher model to generate Triton kernels multiple times, denoted by $m_{i}\in[1,10]$. Each resulting output $o_{i,j}$ (where $j\in\{1,\dots,m_{i}\}$) consists of a synthesized reasoning trace followed by the corresponding Triton implementation. For a given teacher model $\pi_{\text{teacher}}$, this process yields an SFT dataset:
$\mathcal{D}_{SFT}=\big\{(q_{i},o_{i,j})\big\}_{g_{i}\in\mathcal{G}_{KB},q_{i}=q(g_{i}),\{o_{i,j}\}_{j=1}^{m_{i}}\sim\pi_{\text{teacher}}(\cdot|q_{i})}.$
With the SFT dataset, we train our base model parameterized by $\theta$, $\pi_{\theta}$, to maximize the log-likelihood:
$\max_{\theta}\mathbb{E}_{(q_{i},o_{i,j})\sim\mathcal{D}_{SFT}}\left[\log\pi_{\theta}(o_{i,j}|q_{i})\right].$
Through SFT, it learns to internalize the underlying logic of kernel optimization rather than simply memorizing syntax. To explore the impact of different teacher expertise, we construct two separate SFT datasets (approximately 60K samples each) using DeepSeek R1 (Guo et al., 2025) and GPT OSS 120B (OpenAI, 2025) respectively, resulting in two distinct SFT models. Examples of task queries and outputs are provided in Appendix G.5, with additional training details in Appendix E.
2.1.2 Data Augmentation and Difficulty-Aware Data Mixing
The selection of training data is crucial for effective RL post-training, and many works have shown that selective sampling by leveraging additional information, such as difficulty or interaction between data points, can significantly improve model performance (Yu et al., 2025; Chen et al., 2025, 2024). While KernelBook provides a large-scale and diverse set of kernels, naive uniform sampling may lead to suboptimal efficiency. Consequently, we augment each task and apply a difficulty aware mixing strategy to refine the RL training set.
Diverse Input Augmentation.
We start from the KernelBook task set $\mathcal{G}_{KB}=\{g_{i}\}_{i\in\{1,\dots,11K\}}$, used to construct the SFT datasets. To increase diversity and robustness, we augment each task $g_{i}$ with additional input-generating functions using GPT-OSS 120B, producing multiple tasks $\{\tilde{g}_{i,j}\}_{j\in\{1,\dots,n_{i}\}}$ with $n_{i}$ different input shapes. After augmentation, we validate the tasks by execution, so the number of valid tasks $n_{i}\leq 5$ may vary. We denote the resulting set of augmented tasks as $\tilde{\mathcal{G}}_{KB}=\{\tilde{g}_{i,j}\}_{i\in\{1,\dots,|\mathcal{G}_{KB}|\},j\in\{1,\dots,\tilde{m}_{i}\}}$. By default we use the original task set $\mathcal{G}_{KB}$; when RL training uses the augmented set $\tilde{\mathcal{G}}_{KB}$ we indicate this with “+ IA”. In Section 3.2, we analyze the impact of input augmentation on final performance.
Difficulty-Aware Data Mixing.
Then, following the difficulty levels defined in KernelBench (Ouyang et al., 2025), we classify tasks into two levels, where Level 1 represents single-kernel tasks, such as convolution, and Level 2 represents fusion tasks, such as conv+bias+ReLU. To label each task $g_{i}$, we utilize an LLM based labeler (Qwen3-235B-Instruct (Team, 2025)) and let the difficulty label as $d_{i}=d(g_{i})\in\{1,2\}$. Focusing on Level 1 and Level 2 tasks, we denote the subset of tasks at difficulty level $d$ as $\mathcal{G}_{KB}^{(d)}=\{g_{i}\in\mathcal{G}_{KB}|d(g_{i})=d\}$ for $d\in\{1,2\}$ and construct RL training datasets according to various mixture probabilities $\bm{p}$ as follows:
$\mathcal{D}_{RL}(\bm{p})=\{(g_{i},q_{i})\}_{d\sim\bm{p},g_{i}\sim\mathcal{G}_{KB}^{(d)},q_{i}=q(g_{i})},$
where $\bm{p}=(p^{(1)},p^{(2)})\in[0,1]^{2}$ where $p^{(d)}$ denotes the sampling probabilities for difficulty level $d$. In our experiments (Section 3.2), we explore different training data mixtures and evaluate their effectiveness during RL post-training to identify the optimal configuration. Throughout this paper, we set $\bm{p}=(1.0,0.0)$ unless otherwise specified and omit $\bm{p}$ for notational simplicity. Additional details regarding the labeling prompt is provided in Appendix G.1.
Building on the model equipped with essential Triton skills, $\pi_{\theta}$ and a curated RL dataset $\mathcal{D}_{RL}$, we now focus on the core of our RL framework. Before diving into the specifics, we outline the overall RL training workflow of TritonRL. For each task $(g_{i},q_{i})$ sampled from the RL training dataset $\mathcal{D}_{RL}$, the model generates a group of outputs $\{o_{i,j}\}_{j=1}^{G}$ for group size $G$, where each output $o_{i,j}=\{o_{i,j}^{\text{plan}},o_{i,j}^{\text{code}}\}$ consists of reasoning traces planning on how to optimize Triton kernels ($o_{i,j}^{\text{plan}}$) followed by Triton code ($o_{i,j}^{\text{code}}$). After executing the generated Triton code, the model receives reward feedback $r_{i,j}$ based on the quality of the output, which it uses to update its parameters $\theta$ to maximize expected rewards over time. We illustrate the overall RL training workflow of TritonRL in Figure 1.
@@ 2.1.2 Data Augmentation and Difficulty-Aware Data Mixing
Figure 2: Illustration of the flow of our robust verifier incorporating syntax and functionality checkers and the examples of invalid Triton codes. (a) invalid syntax: the code lacks any Triton blocks with @triton.jit. (b) invalid functionality: the code include dummy Triton code that just copies data without meaningful operation delegating core operation (matrix multiplication) to PyTorch modules (torch.matmul).
2.2 Reward Formulation with Robust Verification
In reinforcement learning, reward design is critical, as poorly aligned signals can lead to reward hacking, where the model exploits loopholes rather than achieving the intended objectives. We identify specific Triton failure modes, such as ignoring core kernel functionality, that are often missed by conventional reward designs based solely on unit tests. To mitigate this, we develop fine grained verification layers that rigorously assess code quality across multiple dimensions.
For a generated output $o$ and a corresponding reference PyTorch code $g$, we introduce verifiers that evaluate the following aspects:
  • valid: A binary verifier to check output code is syntactically and functionally valid Triton code. It is defined as $\texttt{valid}(o)=\texttt{syntax}(o)\cdot\texttt{func}(o)$, where:
@@ 2.2 Reward Formulation with Robust Verification
    • func: A binary functionality verifier to detect whether $o^{\text{code}}$ constitutes a valid Triton kernel. Syntax checks alone are insufficient, since models may output code that superficially passes verification but defers computations to high-level PyTorch modules (e.g., torch.nn, @) or hardcodes constants, as in Figure 2 (b). To address this, we combine a rule-based linter, which detects actual calls of Triton kernels and flags reliance on PyTorch modules, combined with an LLM-based judge (Qwen3-235B-Instruct (Team, 2025)) that evaluates semantic correctness against task specifications. In Appendix H, we provide additional details on the LLM judge used for functionality verification.
    See Figure 2 for the illustration of invalid Triton codes that fail syntax and functionality checks.
@@
  • compiled: A binary verifier that checks whether a Triton code can be successfully compiled without errors.
@@
  • correct: A binary verifier that evaluates whether a Triton code produces correct outputs by compiling and comparing its results against those of the reference code $g$ for five random test inputs $\{x_{l}\}_{l=1}^{5}$.
@@
    $\begin{aligned} \texttt{correct}(g,o)\;&=\;\texttt{compiled}\!\left(o\right)\cdot\prod_{l=1}^{5}\mathbbm{1}\!\left[\,o^{\text{code}}(x_{l})==g(x_{l})\,\right]\end{aligned}$
@@
  • speedup: A scalar score that quantifies the execution time improvement of the generated Triton code $o^{\text{code}}$ relative to the reference code $g$. For a test input $x$,
@@
    $\begin{aligned} \texttt{speedup}(g,o)=\frac{\tau(g,x)}{\tau(o^{\text{code}},x)}\cdot\texttt{correct}(g,o),\end{aligned}$
@@
    where $\tau(\cdot,x)$ is the average runtime of a given code for input $x$ measured over ten runs.
Figure 3: Side-by-side comparison of fine-tuned models for Triton programming (KernelLLM and AutoTriton) and their base models. For each metric, the numbers represent the ratio of corresponding samples among 10 generated samples for each of the 100 tasks in KernelBench Level 1. TritonRL (D) denotes our model trained from SFT model distilled from DeepSeek-R1. KernelLLM and AutoTriton show increased functional invalidity (red) after fine-tuning, while TritonRL significantly reduces such errors.
In Figure 3, we analyze the validity of Triton codes generated by existing specialized models, such as AutoTriton (Li et al., 2025b) and KernelLLM (Fisches et al., 2025), and observe a high frequency of functionality violations (red). Interestingly, their base models exhibit lower error rates prior to fine-tuning. This suggests that fine-tuning processes that lack rigorous functionality checks can inadvertently incentivize models to adopt invalid shortcuts, superficially satisfying syntax requirements while failing to implement the core kernel operations.
To mitigate these failure modes, we define two reward functions that deliver robust, accurate signals for correctness and efficiency as follows:
$\begin{aligned} R^{\text{correct}}(g,o)&=\texttt{valid}(o)\cdot\texttt{correct}(g,o),\\ R^{\text{speedup}}(g,o)&=\texttt{valid}(o)\cdot\text{clip}\left(\texttt{speedup}(g,o),2\right).\end{aligned}$
The syntax and functionality checks both serve as necessary conditions for both reward signals to prevent the model from being incentivized to exploit shortcuts. Also, we clip the speedup at $2$ to prevent extreme values from destabilizing training.
2.3 Policy Optimization with Hierarchical Reward Decomposition
Building upon the robust reward designs, we introduce a policy optimization framework designed to leverage these signals effectively. While the verification layers ensure high-fidelity feedback, training language models with long reasoning traces remains challenging due to the credit assignment problem. When a single, monolithic reward is applied to a lengthy response, it fails to distinguish between the tokens that contribute to a brilliant optimization strategy and those that contain a minor syntax error in the implementation.
This issue is particularly pronounced in Triton kernel code generation, where a model may propose a promising memory orchestration plan in its reasoning trace, yet fail the entire task due to a low level indexing error in the subsequent code. Uniformly penalizing such a response conflates high quality reasoning with poor execution (Qu et al., 2025), preventing the model from internalizing effective optimization strategies.
GRPO with Hierarchical Reward Decomposition (HRD).
To address this, we propose an optimization approach based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), augmented with Hierarchical Reward Decomposition (HRD). GRPO is an actor-only RL algorithm that optimizes the policy by comparing the rewards of multiple outputs within a group generated for the same query, eliminating the need for a separate critic model.
Our key innovation is to segment the long output sequence into two hierarchical actions: high-level planning ($o^{\text{plan}}$) and low-level coding ($o^{\text{code}}$). By assigning distinct reward credits to these segments (Figure 1), we better align the Triton kernel optimization plan with the final implementation. We decompose the objective into two components and optimize them jointly: